# Complementary Advantages: Exploiting Cross-Field Frequency Correlation for NIR-Assisted Image Denoising

### v1

@@ Abstract
Existing single-image denoising algorithms often struggle to restore details when dealing with complex noisy images. The introduction of near-infrared (NIR) images offers new possibilities for RGB image denoising. However, due to the inconsistency between NIR and RGB images, the existing works still struggle to balance the contributions of two fields in the process of image fusion. In response to this, in this paper, we develop a cross-field Frequency Correlation Exploiting Network (FCENet) for NIR-assisted image denoising. We first propose the frequency correlation prior based on an in-depth statistical frequency analysis of NIR-RGB image pairs. The prior reveals the complementary correlation of NIR and RGB images in the frequency domain. Leveraging frequency correlation prior, we then establish a frequency learning framework composed of Frequency Dynamic Selection Mechanism (FDSM) and Frequency Exhaustive Fusion Mechanism (FEFM). FDSM dynamically selects complementary information from NIR and RGB images in the frequency domain, and FEFM strengthens the control of common and differential features during the fusion of NIR and RGB features. Extensive experiments on simulated and real data validate that our method outperforms various state-of-the-art methods in terms of image quality and computational efficiency. The code will be released to the public.
@@ 1 Introduction
Due to the inherent physical limitations of digital imaging devices, the captured images are often affected by various types of noise, which have a significant impact on applications such as 24-hour surveillance, autonomous driving, and smartphone photography. Restoring clean detailed information from noisy images that lack substantial information is highly ill-posed. Although the development of deep learning techniques [31, 20] that have brought forth a variety of image denoising methods
Figure 1: Comparisons of PSNR, FLOPs, and Parameters on the DVD dataset [16] are presented. Green circles represent the single-image denoising algorithm, and blue circles represent the NIR-assisted image denoising algorithm. The circle radius represents the number of parameters. Our method (red circles) achieves the superior performance while maintaining efficiency.
[1, 43, 12, 47], existing algorithms still struggle to restore details when dealing with complex noisy images, especially in low-light environment. Therefore, it is necessary to introduce supplementary information from new fields. Fortunately, the introduction of near-infrared (NIR) images offers new possibilities for image denoising. NIR images with high Signal to Noise Ratio (SNR) can be captured at a low cost and is suitable for everyday photography. Thus, NIR-assisted image denoising [42, 22, 32] has become a promising solution for image denoising.
@@
In existing research, Some methods attempt to directly integrate inconsistent features from two fields [39, 38, 26, 27]. However, these methods indiscriminately fuse cross-field information, relying on the fitting capabilities of neural networks to learn the fusion of NIR and RGB images under spatial inconsistency, which often leads to suboptimal denoising results and lower computational efficiency.
Some methods reinforce common features from NIR and RGB images and diminish the weight of inconsistent features during the fusion [42, 17, 10, 16]. These image fusion schemes typically only enhance the common information in NIR and RGB during fusion, but some differential information such as the sharp textures in NIR and the color information in RGB are discarded. In this paper,
@@
Figure 2: Superiority of our method. Visual comparisons on a challenging noisy RGB-NIR image pair. Pay attention to the differences in structural and color information between two fields. Our method produces a better denoising result with clear details and fewer artifacts.
@@
we developed an cross-field Frequency Correlation Exploiting Network (FCENet) for NIR-assisted image denoising. We conduct a series of analyses on the characteristics of NIR and RGB images in the frequency domain and obtain a key observation: in the same scene, the similarity between noisy RGB images and clean RGB images decreases from low to high frequency, whereas the similarity between NIR images and clean RGB images increases from low to high frequency. This significant cross-field frequency correlation prior provide us with inspiration, prompting us to subsequently establish detailed frequency domain learning to thoroughly excavate complementary information from NIR and RGB images.
@@
Specifically, our frequency domain fusion network consists of two mechanisms: The Frequency Dynamic Selection Mechanism (FDSM) and Frequency Exhaustive Fusion Mechanism (FEFM). In order to comprehensively extract useful features from inconsistent images, we formulate the FDSM based on dynamic filters guided by the results observed in the frequency domain. The FDSM efficiently screens complementary features from NIR and RGB in the frequency domain, providing high quality material for subsequent image fusion. Then, to ensure thorough frequency domain fusion, we innovatively formulate FEFM. In FEFM, for the common and differential features in the NIR and RGB feature maps output by FDSM, we simultaneously model local and long-range correlations across fields to fully model the common features from cross-field information, and employ a differential cross-attention approach to complement some key differential high frequency features in NIR. Our contributions are summarized as follows:
@@
  • We establish the frequency correlation prior between NIR and RGB images through a series of frequency domain analyses of two fields, which prompt the implementation of integrating complementary features in the frequency domain.
@@
  • We propose an efficient frequency dynamic selection mechanism on the basis of the frequency correlation prior to identify which useful low-frequency and high frequency information should be retained in NIR and RGB features.
@@
  • Extensive experiments on simulated and real data validate that our method outperforms various state-of-the-art (SOTA) methods in terms of image quality and computational efficiency.
@@ 2.1 Single Image Denoising
Image denoising [30, 19] is a classic task in low-level vision. In recent years, the theory of deep learning has greatly improved the accuracy and efficiency of denoising [2, 1, 36, 6]. DnCNN [46] uses a simple network with batch normalization and residual learning, which outperforms traditional denoising methods. MPRNet [43] designed a multi-stage restoration framework, and Restormer [44] proposed an efficient transformer framework for high-resolution images. Xformer [45] proposed a transformer framework that combines spatial self-attention and channel self-attention. However, despite their strong denoising capabilities, single-image denoising still inevitably loses a lot of spatial information, leading to overly smooth edge textures in the denoised images.
2.2 NIR-assited Image Restoration
To achieve better denoising effects, researchers have attempted to use near-infrared (NIR) images to assist in RGB image denoising [42, 22, 7, 33]. Compared with single-image recovery, NIR images can help restore the details of degraded images [38, 32]. In early research, Yan et al. [42] proposed a Scale Map from the perspective of gradient inconsistency to find universal usable edges and smooth transitions in NIR images. With the widespread application of deep learning methods, CUNet [10] uses a convolutional sparse coding model to extract common features from cross-field images. FGDNet [26] and SANet [27] fuse the features of two fields in the frequency domain. MNNet [39] proposed an observation model that takes into account the modality gap between the target and guiding images. DVN [16] integrates structural inconsistency priors into deep features, specifically strengthening the common deep features in NIR and RGB. NAID [38] proposed an efficient Selective Fusion Module (SFM), which can be plugged into advanced image restoration networks. However, when confronted with dense noise in extreme environments, most existing NIR-assisted image denoising methods struggle to achieve a balance between color restoration, detail preservation, and artifact suppression under inconsistency.
@@ 2.3 Frequency Domain Image Restoration
Frequency analysis in digital signal processing emphasizes the frequency properties of signals [24, 28, 40]. Recently, frequency domain analysis has shown great potential in improving the performance of various learning-based image restoration frameworks [15, 41, 49, 37]. In [18], An efficient estimation of scaled dot-product attention based on a frequency domain self-attention solver (FSAS) was proposed. HDNet [14] proposed a high-resolution dual-domain learning network. SFNet [9] uses filters to dynamically decouple feature mappings into different frequency components and adaptively extract useful features. In this paper, we emphasize the importance of frequency domain information for the research on fusion under RGB/NIR inconsistency and design a comprehensive frequency domain fusion scheme for this task to ensure the utilization of complementary information from NIR and RGB images.
@@ 3 Method
(a) Frequency Correlation between NIR and RGB
(b) Visualization Results after Filtering
Figure 3: The analysis of NIR and RGB frequency correlation. (a) The left graph represents the frequency correlation between noisy RGB and clean RGB and the right graph represents the frequency correlation between near-infrared images and clean RGB, with the horizontal axis indicating the cutoff frequency of the high-pass filter. Different colored curves represent different scenarios. (b) The visualization results of images output by a fixed frequency filter, with the first row showing high frequency (HF) part and the second row showing low frequency (LF) part.
we first introduce the cross-field frequency correlation prior of NIR and RGB images. Based on the cross-field frequency correlation priors, we formulate a detailed frequency domain scheme, including the Frequency Dynamic Selection Mechanism (FDSM) and the Frequency Exhaustive Fusion Mechanism (FEFM) for thorough frequency exploitation.
@@ 3.1 Cross-field Frequency Correlation Prior
To address the fusion of NIR and RGB images under spatial inconsistency, it is imperative to distinguish which information in NIR and RGB images is useful and which is redundant or even erroneous for denoising. We randomly select 30 pairs of NIR images, noisy RGB images, and clean RGB images from the existing RGB-NIR dataset and performed Fourier Transform on them. To facilitate the study of the correlation between NIR and RGB in the frequency domain, we apply a high-pass filter with a fixed cutoff frequency to the aforementioned three kinds of frequency domain images and convert the filtered results back to the spatial domain. We calculate the Structural Similarity (SSIM) [35] between NIR images and noisy RGB images as well as clean RGB images at different cutoff frequencies as a measure of frequency domain correlation $S$, as shown in Fig. 3(a). The process can be described as
$S=\mathrm{SSIM}(\mathcal{F}^{-1}(\mathrm{H}(F_{target})),\mathcal{F}^{-1}(\mathrm{H}(F_{gt}))),$ (1)
where $\mathcal{F}^{-1}$ denote 2D-IDFT operations, $F_{target}$ and $F_{gt}$ represent NIR or noisy RGB and clean RGB after 2D-DFT transformation. $\mathrm{H}$ represents the filtering of the input image with a high-pass filter. We find that: in the main energy frequency, the image similarity between the input noisy RGB and clean RGB images decreases from low to high frequencies, while the image similarity between the input NIR image and clean RGB image increases from low to high frequencies. More NIR-RGB paired data also exhibit the same correlation (see supplementary materials). Fig. 3(b) also clearly shows that for noisy RGB images, their high-frequency information is mainly composed of noise and lacks detail information, while the low-frequency information more faithfully reflects the color information in the original image, and the NIR image is just the opposite. These cross-field Frequency correlation priors indicate that distinguishing useful information from NIR and RGB image inputs in the frequency domain is simple and efficient, which prompts us to propose a comprehensive frequency domain learning framework.
Figure 4: The overall architecture of the proposed cross-field Frequency Correlation Exploiting Network (FCENet).
@@ 3.3 Frequency Dynamic Selection Mechanism
The frequency correlation prior makes us realize that extracting cross-field complementary information from the frequency domain is efficient. For feature extraction, traditional convolutional kernels share weights across the entire feature map, lacking flexibility to address the numerous inconsistencies in NIR and RGB images, which affects the efficiency of extraction. We introduce frequency domain dynamic convolution to adapt to specific inputs and effectively filter out complementary features.
@@
For dynamic filters, some works [48, 23] that generate deep separable and spatially variant dynamic filtering kernels are usually computationally intensive and time-consuming, and require a large amount of data. We employed a method similar to [5, 29] by generating dynamic weights to dynamically aggregate multiple parallel convolutional kernels related to the input, and performing this process in the frequency domain. Specifically, the aggregated features $\mathbf{A_{R}}$ and $\mathbf{A_{N}}$ are used to generate the weights for $k$ learnable frequency domain filter bases. For each feature channel, these filter bases are linearly combined to form a frequency domain global filter, which is used to select useful frequency domain information from the input NIR and RGB features. Finally, $\mathbf{F_{R}}$, $\mathbf{F_{N}}$ denoting the filtered useful features of NIR and RGB can be represented as
$\mathbf{DF_{I}}=\mathrm{Softmax}(\mathrm{MLP}(\mathrm{Pooling}(\mathbf{A_{I}})))\cdot\mathbf{G},$ (2)
$\mathbf{F_{I}}=\mathcal{F}^{-1}(\mathcal{F}(\mathbf{I})\odot\mathbf{DF_{I}}),$ (3)
where $\mathbf{DF_{I}}$ is frequency domain dynamic filter, and I represents $\mathbf{R}$ or $\mathbf{N}$. $\mathbf{G}\in\mathbb{R}^{M\times\ k}$ represents the combination of $k$ learnable frequency domain filtering kernels, and M is the spatial dimension of the convolutional kernel. $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the 2D-DFT and 2D-IDFT operations, respectively.
@@ 3.4 Frequency Exhaustive Fusion Mechanism
After the NIR and RGB feature maps pass through the Frequency Domain Selection Module (FDSM), the useful features are retained. Next, in order to fuse the two inconsistent fields, we further integrate the information of the two fields in the frequency domain under the incentive of previous frequency domain correlations. Specifically, our Frequency Exhaustive Fusion Mechanism (FEFM) includes two key components: Common Feature Reinforcement Mechanism (CRM) and Differential Feature Reinforcement Mechanism (DRM). CRM comprehensively explores the correlation between NIR and RGB features in the frequency domain to strengthen the modeling of common features, while DRM employs differential modeling to enhance key differential high-frequency features.
Common feature Reinforcement Mechanism (CRM). The convolution theorem establishes that the correlation or convolution of two signals in the spatial domain is equal to their Hadamard product in the frequency domain. Leveraging this property, previous work [18] integrate the frequency domain into the self-attention mechanism, simplifying matrix multiplication into lightweight element-wise dot product operations. Although the dot product is efficient, it only calculates the point-wise correlation coefficients in the frequency domain, missing the long-range correlations in the frequency domain. To more comprehensively model the common features of NIR and RGB, We apply a point-wise convolution and a depthwise convolution to the input NIR and RGB feature maps $\mathbf{F_{R}}$, $\mathbf{F_{N}}$, to obtain encoded features $\mathbf{Q}{=}W_{d}^{Q}W_{p}^{Q}\mathbf{F_{R}}$, $\mathbf{K}{=}W_{d}^{K}W_{p}^{K}\mathbf{F_{N}}$, $\mathbf{V}{=}W_{d}^{V}W_{p}^{V}\mathbf{F_{N}}$. To facilitate the calculation of frequency domain correlations, we apply 2D-DFT to the input features to produce frequency cubes $\mathbf{D_{Q}}$ and $\mathbf{D_{K}}$. We simultaneously modeled the long-range correlations and point-wise correlation coefficients of NIR and RGB features in the frequency domain, resulting in a correlation map that integrates local similarities and long-range channel correlations within the frequency domain. The output of CRM can be represented as
$\mathbf{O_{CRM}}=(\mathbf{D_{Q}}\odot\mathbf{D_{K}})\cdot\mathrm{Softmax}(\mathbf{D_{Q}}\cdot\mathbf{D_{K}^{\top}}/\alpha),$ (4)
where $\alpha$ is a learnable scaling parameter.
@@
Uformer (CVPR 2022) 31.34 0.949 29.12 0.927 27.58 0.908 9.60 20.63
Restormer (CVPR 2021) 31.30 0.949 29.20 0.928 27.76 0.910 36.06 26.13
MPRNet (CVPR 2021) 31.78 0.951 29.37 0.929 27.84 0.910 133.51 15.74
Xformer (ICLR 2024) 31.41 0.950 29.30 0.928 27.43 0.908 39.34 25.22
CUNet (TPAMI 2020) 28.76 0.923 26.81 0.898 25.64 0.875 6.72 0.44
SANet (CVPR 2023) 30.13 0.938 27.84 0.917 26.41 0.901 25.01 5.38
MNNet (IF 2022) 30.14 0.944 28.26 0.923 26.75 0.906 10.97 0.76
DVN (AAAI 2022) 31.50 0.955 29.62 0.940 28.26 0.927 48.30 6.97
NIR-Restormer (arXiv 2024) 31.81 0.957 29.81 0.942 28.31 0.928 45.31 30.26
@@
FCENet (Ours) 32.43 0.963 30.37 0.950 28.78 0.939 19.52 4.32
Differential feature Reinforcement Mechanism (DRM). Traditional cross-attention [21, 34, 13] structures are typically used to capture common features but do not effectively utilize differential information. Although this approach is correct in general image fusion frameworks, in this paper, the features of the two fields are considered useful after passing through the frequency domain selection module. By observing the feature maps (Fig. 9), we find that the high-frequency texture information in the NIR features is clear compared to the RGB feature maps. The common features obtained by weighted fusion through cross-attention remain somewhat blurry, which greatly affects the denoising effect. These differential high-frequency features on NIR are crucial for image restoration. We adopt a differential cross-attention to complement the missing differential information, and the output of CRM can be represented as
$\begin{split}\mathbf{O_{DRM}}=(\mathbf{V}-\lambda\ \mathbf{V}\odot\mathcal{F}^{-1}(\mathbf{O_{CRM}})),\end{split}$ (5)
where $\lambda$ is a learnable scalar that dynamically controls the weights of two feature maps. Although DRM effectively models the differential high-frequency information, the fused features lack a significant amount of common features and background information. To further refine the fused features, we perform a dot product between the encoded features Q and the output of CRM, and add the result to the output of DRM as the final fused result.

### v2

@@ Abstract
Existing single-image denoising algorithms often struggle to restore details when dealing with complex noisy images. The introduction of near-infrared (NIR) images offers new possibilities for RGB image denoising. However, due to the inconsistency between NIR and RGB images, the existing works still struggle to balance the contributions of two fields in the process of image fusion. In response to this, in this paper, we develop a cross-field Frequency Correlation Exploiting Network (FCENet) for NIR-assisted image denoising. We first propose the frequency correlation prior based on an in-depth statistical frequency analysis of NIR-RGB image pairs. The prior reveals the complementary correlation of NIR and RGB images in the frequency domain. Leveraging frequency correlation prior, we then establish a frequency learning framework composed of Frequency Dynamic Selection Mechanism (FDSM) and Frequency Exhaustive Fusion Mechanism (FEFM). FDSM dynamically selects complementary information from NIR and RGB images in the frequency domain, and FEFM strengthens the control of common and differential features during the fusion process of NIR and RGB features. Extensive experiments on simulated and real data validate that the proposed method outperforms other state-of-the-art methods. The code will be released at https://github.com/yuchenwang815/FCENet.
@@ 1 Introduction
Due to the inherent physical limitations of digital imaging devices, the captured images are often affected by various types of noise, which have a considerable impact on applications such as 24-hour surveillance, autonomous driving, and smartphone photography. Although the development of deep learning techniques [36, 24] has brought forth a variety of image denoising methods
Figure 1: Comparisons of PSNR, FLOPs, and Parameters on the DVD dataset [18] are presented. Green circles represent the single-image denoising algorithms, and blue circles represent the NIR-assisted image denoising algorithms. The circle radius represents the number of parameters. The proposed method (red circles) achieves the superior performance while maintaining efficiency.
[1, 47, 13, 51, 3, 28], existing algorithms still find it difficult to deal with complex noisy images, especially in low-light environment. Restoring clean detailed information from noisy images that lack substantial information is highly ill-posed. Therefore, it is necessary to introduce supplementary information from new fields. Fortunately, NIR images with high Signal to Noise Ratio (SNR) can be captured at a low cost and are suitable for everyday photography. Thus, NIR-assisted image denoising [46, 26, 37] has become a promising solution for image denoising.
@@
In existing researches, some methods attempt to directly integrate inconsistent features from two fields [44, 43, 31, 32]. However, these methods indiscriminately fuse cross-field information, relying on the fitting capabilities of neural networks to learn the fusion of NIR and RGB images under spatial inconsistency, which often leads to suboptimal denoising results and lower computational efficiency. Some methods reinforce common features from NIR and RGB images and diminish the weight of inconsistent features during the fusion [46, 19, 11, 18]. These image fusion schemes typically only enhance the common information in NIR and RGB during fusion, but some differential information such as the sharp textures in NIR and the color information in RGB is discarded.
In this paper,
@@
Figure 2: Visual comparisons on the challenging noisy RGB-NIR image pairs. Pay attention to the differences in structural and color information between two fields. Our method produces a better denoising result with clear details and fewer artifacts.
@@
we develop a cross-field Frequency Correlation Exploiting Network (FCENet) for NIR-assisted image denoising. We conduct a series of analyses on the characteristics of NIR and RGB images in the frequency domain and obtain a key observation: in the same scene, the similarity between noisy RGB images and clean RGB (noise-free RGB) images decreases from low to high frequency, whereas the similarity between NIR images and clean RGB images increases from low to high frequency. This significant cross-field frequency correlation prior redefines cross-field fusion as a process of frequency domain adaptive selection, motivating us to subsequently establish detailed frequency domain learning to precisely control the restoration of texture details and colors.
@@
Specifically, the frequency domain fusion network consists of two mechanisms: Frequency Dynamic Selection Mechanism (FDSM) and Frequency Exhaustive Fusion Mechanism (FEFM). In order to comprehensively extract useful features from inconsistent images, the FDSM is formulated based on dynamic filters under the guidance of frequency domain prior. The FDSM efficiently screens complementary features from NIR and RGB in the frequency domain, providing high quality material for subsequent image fusion. Then, the FEFM is formulated to ensure thorough frequency domain fusion. The FEFM integrates local and long-range correlations across fields to fully extract the common feature from cross-field information, and employs a differential cross-attention approach to complement some key differential high frequency features in NIR features.
The proposed FCENet can utilize complementary information from cross-field images. Extensive experiments on simulated and real data validate that our method outperforms other state-of-the-art methods in terms of image quality and computational efficiency. The main contributions of this work are summarized as follows:
@@
  • We establish the frequency correlation prior between NIR and RGB images through a series of frequency domain analyses of two fields, which prompts the integration of complementary features in the frequency domain.
@@
  • We propose an efficient frequency dynamic selection mechanism based on the frequency correlation prior to identify which useful low-frequency and high frequency information should be retained in NIR and RGB features.
@@ 2.1 Single Image Denoising
Image denoising [29, 12] is a classic task in low-level vision. In recent years, deep learning has greatly improved the accuracy and efficiency of denoising [2, 7, 35, 22, 14, 23]. DnCNN [50] uses a simple network with batch normalization and residual learning, which outperforms traditional denoising methods. MPRNet [47] designs a multi-stage restoration framework, and Restormer [48] proposes an efficient transformer framework for high-resolution images. Xformer [49] proposes a transformer framework that combines spatial self-attention and channel self-attention. However, despite their strong denoising capabilities, single-image denoising still inevitably loses a lot of spatial information, leading to overly smooth edge textures in the denoised images.
2.2 NIR-Assisted Image Restoration
To achieve better denoising effects, researchers have attempted to use NIR images to assist in RGB image denoising [46, 26, 8, 38]. Compared with single-image recovery, NIR images can help restore the details of degraded images [43, 37]. In early research, Yan et al. [46] proposes a scale map from the perspective of gradient inconsistency to find universal usable edges and smooth transitions in NIR images. With the proliferation of deep learning, CUNet [11] uses a convolutional sparse coding model to extract common features from cross-field images. FGDNet [31] and SANet [32] fuse the features of two fields in the frequency domain. MNNet [44] proposes an observation model that takes into account the modality gap between the target and guiding images. DVN [18] integrates structural inconsistency priors into deep features, specifically strengthening the common deep features in NIR and RGB images. NAID [43] proposes Selective Fusion Module (SFM), which can be plugged into advanced image restoration networks. However, under extreme noisy conditions, existing NIR-assisted denoising methods fail to balance color accuracy, detail retention, and artifact suppression.
@@ 2.3 Frequency Domain Image Restoration
Frequency analysis in digital signal processing emphasizes the frequency properties of signals [33, 45]. Recently, frequency domain analysis has shown great potential in improving the performance of various learning-based image restoration frameworks [17, 53, 42]. In [21], An efficient estimation of scaled dot-product attention based on a frequency domain self-attention solver (FSAS) is proposed. HDNet [16] proposes a high-resolution dual-domain learning network. SFNet [10] uses filters to dynamically decouple feature mappings into different frequency components and adaptively extracts useful features. In this paper, we emphasize the importance of frequency domain information for the research on fusion under inconsistency and design a comprehensive frequency domain fusion scheme for this task to ensure the utilization of complementary information from NIR and RGB images.
@@ 3 Method
Figure 3: The analysis of NIR and RGB frequency correlation. (a) The frequency correlation between noisy RGB and clean RGB, with the horizontal axis indicating the cutoff frequency of the high-pass filter. Different colored curves represent different scenarios. (b) The frequency correlation between NIR and clean RGB. (c) The visualization results of images output by a fixed frequency filter, with the first row showing high frequency (HF) part and the second row showing low frequency (LF) part.
we first introduce the cross-field frequency correlation prior of NIR and RGB images. Based on the cross-field frequency correlation prior, a detailed frequency domain scheme is formulated, including the Frequency Dynamic Selection Mechanism (FDSM) and the Frequency Exhaustive Fusion Mechanism (FEFM) for thorough frequency exploitation.
@@ 3.1 Cross-Field Frequency Correlation Prior
To address the fusion of NIR and RGB images under spatial inconsistency, it is imperative to distinguish which information in NIR and RGB images is useful and which is redundant or even erroneous for denoising. We randomly select 30 pairs of NIR images, noisy RGB images, and clean RGB images from the existing RGB-NIR dataset and perform fourier transform on them. To facilitate the study of the correlation between NIR and RGB in the frequency domain, we apply a high-pass filter to the aforementioned three kinds of frequency domain images and convert the filtered results back to the spatial domain. We separately computed the Structural Similarity (SSIM) [40] between NIR/noisy RGB images and clean RGB images at different cutoff frequencies as a measure of frequency correlation $S$, as shown in Fig. 3(a) and Fig. 3(b). The process can be described as
$S=\mathrm{SSIM}(\mathcal{F}^{-1}(\mathrm{H}(F_{target})),\mathcal{F}^{-1}(\mathrm{H}(F_{gt})))$ (1)
where $\mathcal{F}^{-1}(\cdot)$ denote 2D-IDFT operations, $F_{target}$ and $F_{gt}$ represent NIR or noisy RGB and clean RGB after 2D-DFT transformation. $\mathrm{H}(\cdot)$ represents the filtering of the input image with a high-pass filter. The analysis results indicate that, in the main energy frequency, the image similarity between the input noisy RGB and clean RGB images decreases progressively as the frequency increases, while the image similarity between the input NIR image and clean RGB image increases from low to high frequency. Fig. 3(c) also shows that for noisy RGB images, their high frequency information is mainly composed of noise and lacks detail information, while the low frequency information more faithfully reflects the color in the original image, and NIR images are just the opposite. These cross-field Frequency correlation priors indicate that it is simple and efficient to distinguish useful information from NIR and RGB images in the frequency domain, which inspires us to propose a comprehensive frequency domain learning framework.
Figure 4: The overall architecture of the proposed cross-field Frequency Correlation Exploiting Network (FCENet). In the first stage, pre-denoising is performed. In the second stage, NIR image is introduced for NIR-assisted image denoising.
@@ 3.3 Frequency Dynamic Selection Mechanism
The frequency correlation prior indicates that extracting cross-field complementary information from the frequency domain is efficient. For feature extraction, traditional convolutional kernels share weights across the entire feature map, lacking flexibility to address the numerous inconsistencies between NIR and RGB images, which affects the efficiency of extraction. We introduce frequency domain dynamic convolution to adapt to specific inputs and effectively extract complementary features.
@@
For dynamic filters, some works [52, 27] that generate deep separable and spatially variant dynamic filtering kernels are usually computationally intensive and time-consuming, and require a large amount of data. We employ a method similar to [6, 34] by generating dynamic weights to dynamically aggregate multiple parallel convolutional kernels related to the input, and performing this process in the frequency domain. Specifically, the aggregated features $\mathbf{A_{R}}$ and $\mathbf{A_{N}}$ are used to generate the weights for $k$ learnable frequency domain filters. For each feature channel, these filters are linearly combined to form a frequency domain global filter, which is used to select useful frequency domain information from the input NIR and RGB features. Finally, $\mathbf{F_{R}}$, $\mathbf{F_{N}}$ denoting the filtered useful features of NIR and RGB can be represented as
$\mathbf{DF_{I}}=\mathrm{Softmax}(\mathrm{MLP}(\mathrm{Pooling}(\mathbf{A_{I}})))\cdot\mathbf{G}$ (2)
$\mathbf{F_{I}}=\mathcal{F}^{-1}(\mathcal{F}(\mathbf{I})\odot\mathbf{DF_{I}})$ (3)
where $\mathbf{DF_{I}}$ is frequency domain dynamic filter, and $\mathbf{I}$ represents $\mathbf{R}$ or $\mathbf{N}$. $\mathbf{G}\in\mathbb{R}^{M\times\ k}$ represents the combination of $k$ learnable frequency domain filtering kernels, and M is the spatial dimension of the convolutional kernel. $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the 2D-DFT and 2D-IDFT operations, respectively.
@@ 3.4 Frequency Exhaustive Fusion Mechanism
After the NIR and RGB feature maps pass through the FDSM, the useful features are retained. Next, to fuse the two inconsistent fields, we further integrate their information in the frequency domain and propose the FEFM, with this process driven by prior frequency correlations. Specifically, the proposed FEFM includes two key components: Common Feature Reinforcement (CFR) mechanism and Differential Feature Reinforcement (DFR) mechanism. CFR comprehensively explores the correlation between NIR and RGB features in the frequency domain to strengthen the modeling of common features, while DFR employs differential modeling to enhance key differential high frequency features.
Common Feature Reinforcement Mechanism. The convolution theorem establishes that the correlation or convolution of two signals in the spatial domain is equal to their element-wise product in the frequency domain. Leveraging this property, previous work [21] integrate the frequency domain into the self-attention mechanism, simplifying matrix multiplication into lightweight element-wise dot product operations. Although the dot product is efficient, it only calculates the point-wise correlation coefficients in the frequency domain, missing the long-range correlations in the frequency domain. To more comprehensively model the common features of NIR and RGB, we apply a point-wise convolution and a depthwise convolution to the input NIR and RGB feature maps $\mathbf{F_{R}}$, $\mathbf{F_{N}}$, to obtain encoded features $\mathbf{Q}{=}W_{d}^{Q}W_{p}^{Q}\mathbf{F_{R}}$, $\mathbf{K}{=}W_{d}^{K}W_{p}^{K}\mathbf{F_{N}}$, $\mathbf{V}{=}W_{d}^{V}W_{p}^{V}\mathbf{F_{N}}$. To facilitate the calculation of frequency domain correlations, we apply 2D-DFT to the input features to produce frequency cubes $\mathbf{F_{Q}}$ and $\mathbf{F_{K}}$. We simultaneously model the long-range correlation and point-wise correlation of NIR and RGB features in the frequency domain, resulting in a correlation map that integrates local similarity and long-range channel correlation within the frequency domain. The output of CFR can be represented as
$\mathbf{F_{CFR}}=(\mathbf{F_{Q}}\odot\mathbf{F_{K}})\cdot\mathrm{Softmax}(\mathbf{F_{Q}}\cdot\mathbf{F_{K}^{\top}}/\alpha)$ (4)
where $\alpha$ is a learnable scaling
Noisy RGB NIR Restormer MNNet DVN NIR-Restormer Ours Reference
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\sigma=2$ PSNR 31.44 31.50 33.65 32.63 34.05 GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\sigma=4$ PSNR 27.02 27.41 27.28 27.33 28.18 GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\sigma=6$ PSNR 26.37 27.02 27.55 28.54 29.93 GT
@@
Uformer [41] 31.34 0.949 29.12 0.927 27.58 0.908 9.60 20.63
Restormer [48] 31.30 0.949 29.20 0.928 27.76 0.910 36.06 26.13
MPRNet [47] 31.78 0.951 29.37 0.929 27.84 0.910 133.51 15.74
Xformer [49] 31.41 0.950 29.30 0.928 27.43 0.908 39.34 25.22
CUNet [11] 28.76 0.923 26.81 0.898 25.64 0.875 6.72 0.44
SANet [32] 30.13 0.938 27.84 0.917 26.41 0.901 25.01 5.38
MNNet [44] 30.14 0.944 28.26 0.923 26.75 0.906 10.97 0.76
DVN [18] 31.50 0.955 29.62 0.940 28.26 0.927 48.30 6.97
NIR-Restormer [43] 31.81 0.957 29.81 0.942 28.31 0.928 45.31 30.26
@@
FCENet (Ours) 32.43 0.963 30.37 0.950 28.78 0.939 19.52 4.44
parameter.
Differential Feature Reinforcement Mechanism. Traditional cross-attention [25, 39, 15] structures are typically used to capture common features but do not effectively utilize differential information. Although this approach is correct in general image fusion frameworks, in this paper, the features of the two fields are considered useful after passing through the frequency domain selection mechanism. By observing the feature maps (Fig. 9), we find that the high frequency texture information in the NIR features is clear compared to the RGB feature maps. The common features obtained by weighted fusion through cross-attention remain somewhat blurry, which greatly affects the denoising effect. These differential high frequency features on NIR are crucial for image restoration. We adopt a differential cross-attention to complement the missing differential information, and the output of CFR can be represented as
$\begin{split}\mathbf{F_{DFR}}=(\mathbf{V}-\lambda\ \mathbf{V}\odot\mathcal{F}^{-1}(\mathbf{F_{CFR}}))\end{split}$ (5)
where $\lambda$ is a learnable scalar that dynamically controls the weights of two feature maps. Although DFR effectively models the differential high frequency information, the fused features lack some common features and background information. To further refine the fused features, we perform a dot product between the encoded features Q and the output of CFR, and add the result to the output of DFR as the final fused result.